Fermi National Laboratory
 

Operation Summary for July 19, 2010 - July 23, 2010

 

Fermilab Shutdown Report
7/19/10 – 7/23/10

LINAC

Shutdown Week 1 – On of Off Schedule

  • Preaccelerator
    • I- heater supply upgrade 25%
    • I- generator repair 50%
    • Both ion gauge replacements 25%
    • I- source cleaning 25%
  • Electrical work
    • LERF breaker panel replacement 50%
    • LERF water skid wall disconnects 100%
  • Low Energy
    • LRF1 tuner/fingerstock repair 100%
    • LRF1 PA swap 25%
    • Tanks 1 & 2 ion pump replacement 100%
    • Roughing pump maintenance 75%
    • QPS card upgrade 40%
    • Modulator maintenance 50%
    • Buncher RF coupler install 100%
    • LRF4 filament PLC upgrade 0%
    • LRF5 PA filament leak repair 0%
    • Modulator Accopian PS replacement 0%
  • High Energy
    • Solenoid flow meter upgrade 100%
    • KRF1 120V outlet install 0%
    • Distribution skid HX cleaning 50%
    • Debuncher skid pump relocation 100%
    • Debuncher skid controls upgrade 33%
    • Transformer oil processing 20%
    • QPS maintenance 0%
    • CS/PFN maintenance 25%
    • SC tank 3 TLM install 0%
  • Controls work
    • Pulse shifter chassis move 100%
    • LERF ACNET console installation 100%

Booster

Shutdown Week 1 – On Schedule

  • Directional boring for new sump lines
    • One new line has been pulled into the tunnel
    • Preparing for next directional bore
  • BRF1 repaired, inside tuner bar replaced
  • BRF18 tuner replacement progressing
  • BRF3 at MI60 for testing
  • Preparing for cable pulls starting today

Antiproton Source

No written report, on schedule

Main Injector

No written report, on schedule

Recycler

  • Vertical Flying wire replacement (50%)
    • Flying wire installed and vacuum leak tight
    • Bake to start today and continue over the weekend
  • Replace beam valve solenoids (5%)
    • 1/21(?) done so far
  • Recycler vacuum maintenance (not started)
    • Leak check vacuum in 800 line
      • Repair if needed
    • TSPs sublimation
  • Other
    • Power supplies maintenance
    • Controls nodes
    • Repair RF amplifier on Station 3 - Started
    • Install shielding near BV308F

ECOOL jobs update

  • Inside Pelletron tank
    • Electrical work (80%)
      • Regulation circuitry replaced & tested
      • Upgraded modulator installed & tested
      • Replace SPD05 solenoid power supply – Next week
    • Mechanical work (?% - See next slide)
      • Preliminary cleaning and inspection complete
      • Flow switch installation on collector water system – Underway
  • Outside Pelletron tank
    • Lots of ‘small’ maintenance tasks
      • SF 6 Dryer/Recirculation system, SF 6 transfer skid, Compressor, Water chiller – Underway
    • Back flush SPC80 cooling solenoid
      • In the MI tunnel – Does not need to happen during the first 2 weeks of the shutdown

Charging efficiency degradation investigation

  • Cause is not known
    • Various measurements are being
      made to try to locate the origin
      of the problem
      • e.g.: Sheaves alignment
        top to bottom
  • Depending on the findings
    corrective actions may be needed
      • e.g.: replace and/or align
        the sheaves

Tevatron

NuMI

Week 1

  • Re-cored to supervised access, removed shielding wall & installed crane electronics
  • Took water samples from drainage system (18 samples in total)
  • Complete rad surveys in target pile above the target and horn 1 to track decay for future NoVA & LBNE operations
  • Changed target pile pre-filters
  • Install target hall camera system
  • Installed the coffin transport system in shaft & Target hall area
  • Moved NT-02, PH1-01, & PH2-01 out of the morgue to temp storage in t-hall
  • Move PH1-01, & PH2-01 to C0 storage facility
  • Remove electrical service from target pile fan
  • Remove TP sensors & instrumentation
  • Complete survey of muon alcove 4 & install stand for future detector
  • Start electrical upgrade for ANU/NoVA in SB power supply room

Week 2

  • Move NT-02 to C0 storage facility
  • Continue rad surveys above the target & horn 1 in the target pile
  • Move PEI units A & B from electrical support room to MI62 for repairs
  • Inspect Gollum’s Cave for ground water
  • Install & removal of scaffolding for target pile fan repair
  • Remove target pile fan & dehumidification duct work
  • Start removal of target pile fan and motor
  • Start pre-target & decay pipe drain cleaning
  • Start multiwire replacement
  • Removed PM112 & modified stand (conduit interference w/ new install)
  • Removed PM 108, modified stand, & install new MW
  • Complete decay pipe helium instrumentation upgrades
  • Start change out of regeneration motors on dehumidifiers
  • Continue electrical service upgrade for ANU/NoVA
  • Clean target pile chiller & re-lamp target hall (FESS)

Preview of week 3

  • Continue with target pile repair installing new fan and motor
  • Install additional water protection over Absorber and hadron monitor
  • Complete dehumidifier regeneration motor replacement
  • Finish install of PW 112 & continue replacement of multiwires PW 105 & 107
  • Start installation for MINOS chiller & underground cooling loop upgrade
  • Start AH dehumidifier repair of heater controllers

Cryogenic System

Operations/Maintenance Schedule

  • Tevatron at LHe temperatures except as noted above.
  • MSS power outage went well. CHL NRL plant off and we are relying on Linde deliveries (4-5 truckloads per day) to CHL’s 150,000 gallons of tank storage.
  • Continued ringwide vacuum pump maintenance, repairs, and pumpdown of various components to assure stable Tevatron cryogenic system operations.
  • 8 wet engine cold ends were overhauled (8-10 hrs per unit), 1 overhaul in progress today.
  • Ongoing numerous FRIG/Compr controls/instrumentation maintenance/repairs.

MiniBooNE

No written report, on schedule

SY120

  • MSEPs were transported from MI-65 to Enclosure B
  • Two MSEPs were installed, alignment next week

Collider Detector at Fermilab

No written report, on schedule

D-Zero

No written report, on schedule

EE Support

No written report, on schedule

Controls

Past Week:

  • Recovery from Power Outage
  • Computer Room cooling system repair recovery
  • SDA and Save/Restore Databases OS Upgrades
  • MI-39 CAMAC & Timing Links Installed & Operational
  • MAC Room & Computer Room UPS Maintenance
  • General Controls Maintenance & Support

Upcoming Week:

  • Saturday - ACNET Down (7-10 hrs.)
    • Cross Gallery power outage
  • Monday morning - Applications and Device Database OS Upgrade
  • Monday afternoon – ACNET Console Restarts
  • Tuesday morning – Java Engine reboots
  • Tuesday afternoon – MI network maintenance
    • No network communications to MI/RR, NuMI, MB
  • General Controls Maintenance & Support

Alignment

On schedule

FESS

ELECTRICAL & HIGH VOLTAGE

  • MSS – Buss duct cleaning WIP – No findings
  • Minor re-lamping work [call ins/scheduled]
  • SY Enclosure B – increased thermal heater size
  • Tev Ring pond pit PM/inspections WIP
  • Booster sump pump replacement TBS
  • Tev Ring tunnel sump pump E-3 to determine fault

MECHANICAL

  • CDF – Worthington LCW pumps seal leaks {worn out] – 1 to be rebuilt & 1 new on order
  • MI - basket strainers cleaned
  • MI – sump pumps PM completed[MI 8, 30, 60]
  • SY Enclosures – sump pumps PM completed [E & F1 completed]
  • TEV Pond pump intake screens – major Zebra Mussel infestations – Basket strainers cleaning next week

HVAC

  • Linac Pre accelerator – Ed Pac [L074] unit completely inspected – Freon leak repaired – need to repair humidity control
  • Tev ring – Marvairs WIP
  • MI – Marveairs completed last week – will re-check during the 4 th week
  • MI – dehumidifiers/tunnel to start next week
  • MI 31 – Chiller PM completed[rad area]. Trane unit PM WIP – controller/economizer repair[need more discussion on operational requirements]
  • MI 12 – Temperature discharge within specs for the 2 Marveairs
  • MI 65 – Multi stack cooling water filter/intake cleaning/flushing WIP[dirty]

CUB

  • Chillers/Process - #3[Linac 55 LCW] & #4 – cleaning & inspection WIP – Epoxy coatings scheduled next week
  • Booster Heat Exchanger HE-2 cleaning WIP – HE-1 on line
  • PBar Heat Exchanger HE-9 cleaning WIP – HE-8 [tube & shell] on line
  • Booster LCW #2 pump motor burned up – replacement WIP
  • DDC/METASYS server replacement program pre-check WIP[scheduled for CUB shutdown 7/31 –

8/1

OPERATIONS ENGINEERING

  • Booster Gallery West – Dual Temp piping repairs WIP/Fri completion
  • Booster Gallery East – Dual Temp repairs to start Mon 7/26.
  • Booster Sump Pump piping replacement – Directional bore WIP @ east booster parking lot
  • ICW leak @ AP O/MI – WIP – replaced ~20 ft section under two 480 V duct banks.
  • MI 10 – 40/10 HP pond pump 480V electrical feeder replacement to start Mon 7/26

OTHER

  • SAT 7/24 WORK - X Gallery fan/pump room – chilled water piping repair
  • P-3 hot water pump electrical feeder replacement during LINAC power SD

Safety

Continue using the Take Five principles before, during, and after work activities to:

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/23/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory